Skip to content. | Skip to navigation

Personal tools

Troubleshooting

If the instrument does not appear to be functioning properly, the first step is to go through the standard Verification tests. This may immediately isolate the source of the problem.

Known Problems and Solutions


Filter Wheel: The filter detent switch very infrequently fails. This may cause the filter wheel to fail to completely move into position. This problems manifests as a sharp edge with no flux at the bottom or top of an image. The solution is to 'jiggle' the filter wheel slightly using commands on the Camera-Control GUI to get it into position. From the File pulldown menu, select Filter1 or Filter2 and Move+ until the filter is centered. This problem appears to only occur right after the power has been cycled on the electronics. If it is working, it is unlikely to fail.

  • Detector Noise: The detector occassionally exhibits some 60 Hz and higher frequency noise. This problem has been solved several times now. Please let us know if it returns again.


  • Dome flat fields do not work well: Twilight flats are strongly recommended instead of dome flats and should be sufficient for 2% photometry. Note that you will need darks in addition to the twilight flats.


  • There is no dark filter: The dark filter was removed in July 2004. The current procedure to obtain a dark is to place the J and H2 filters in the beam at the same time, as the bandpasses of these filters are mutually exclusive.


  • Distortion: PANIC has a radial distortion that shifts object positions at the corners by approximately 1.8% of the center-to-corner distance. This distortion has been mapped out with a ZEMAX model of the optical design and an IRAF script to correct for the distortion distcor is included in the panic IRAF package.




Document Actions